# MemeIntel: Explainable Detection of Propagandistic and Hateful Memes

### v1

@@ 4.3 Basic Statistics
Table 3 presents the basic statistics for both datasets. The average explanation length is 94 words for Arabic and 85 words for English. Notably, we instructed GPT-4o to generate explanations with fewer than 100 words. Based on manual evaluation (Table 4), we conclude that both the quality and length of the explanations are appropriate.
@@ 5.1 Instructions Dataset
Our approach follows the standard pipeline for aligning LLMs with user intentions and specific tasks through fine-tuning on representative data Zhang et al. (2023); Kmainasi et al. (2024). This process typically involves curating and constructing instruction datasets that guide the model’s behavior, ensuring it generates responses that align with the desired objectives. For our study, the responses include label and explanation. Hence, we created instruction format for both datasets. For the ArMeme dataset, we replicated the experiments for both Arabic and English explanations.
@@ 5.2 Model Selection
As shown in Figure 1, our first experimental phase involves model selection among several recent VLMs, including Llama-3.2 (11b) Dubey et al. (2024), Paligemma 2 (3b) Steiner et al. (2024), Qwen2-vl Wang et al. (2024), and Pixtral (12b) Agrawal et al. (2024).
We evaluate the base models in a zero-shot setting and fine-tune them using an instruction-following paradigm. The instructions prompt the model to generate responses in the format ‘‘Label: (class_label)’’. We use and a regex-based function to extract the predicted labels.
@@ 5.3 Multi-Stage (MS) Optimization Procedure
To emphasize our novel contribution, we introduce a dedicated optimization procedure to train VLM with MemeXplain, which decouples the classification and explanation generation tasks. This approach is designed to first endow the model with strong task-specific representations through classification-only fine-tuning, and then refine its ability to generate coherent, natural explanations.
Stage 1: Classification Fine-Tuning
In this stage, the model is fine-tuned solely on the classification task. The training objective is restricted to predicting the correct class label. This focused objective encourages the model to develop robust, task-specific representations. We use the QLoRA setup described later with a learning rate of $2\times 10^{-4}$ to optimize the model.
Stage 2: Explanation Enhancement
In this stage, the model is further fine-tuned on a combined label-with-explanation dataset. Here, we employ a reduced learning rate ($1\times 10^{-5}$) to gently adapt the model’s parameters for generating the explanations while preserving the classification performance achieved in Stage 1.
To validate the effectiveness of the multi-stage procedure, we compare it against a single-stage (SS) fine-tuning baseline where the model is directly trained on the label-with-explanation dataset. Our ablation studies (detailed in Section 6) demonstrate that the proposed multi-stage approach significantly outperforms the single-stage strategy.
@@
Qwen2 (7b) Base 63.1 0.550 0.242
Qwen2 (7b) FT 27.0 0.149 0.195
@@
Qwen2 (7b) Base 66.4 0.669 0.442
Qwen2 (7b) FT 77.9 0.773 0.753
@@ 5.4 Training Setup
Our fine-tuning experiments utilize QLoRA Dettmers et al. (2023), which combines INT4 quantization with parameter-efficient fine-tuning through Low-Rank Adaptation (LoRA) Hu et al. (2022). In our setup, the base model is quantized to 4-bit precision, with LoRA updates applied to a subset of the model parameters. This approach was selected to address computational resource constraints. Furthermore, deploying models for inference incurs significant costs. Therefore, we focus on quantized models and assessing their performance accordingly.
For all experiments, we fine-tuned the models using the QLoRA approach with 4-bit quantization. This approach was chosen due to its efficiency in reducing memory usage while maintaining model performance. We adapted all relevant submodules (vision, language, attention, and MLP layers) with a LoRA rank of 16, an alpha of 16, and no dropout. For training, we used a per-device batch size of 2 with gradient accumulation over 4 steps and optimized using AdamW with a learning rate of $2\times 10^{-4}$, a weight decay of 0.01, and a linear scheduler with 5 warmup steps. For the second stage experiments (label-with-explanation), the learning rate was reduced to $1\times 10^{-5}$.
@@ 5.5 Evaluation Setup and Metrics
We train the models using the training set, fine-tune the parameters with the development set, and evaluate their performance on the test set as reported in Tables 5 and 6. For performance measurement across different experimental settings, we compute accuracy, weighted F1 score, and macro-F1 score. We evaluate the model’s explanation performance on the test set using semantic similarity-based metric, measured by the F1 score within BERTScore (Zhang et al., 2020). This score is computed using contextual embeddings extracted from pre-trained BERT models. To enhance accuracy, we utilize language-specific transformer models for embedding extraction. For Arabic we use AraBERT (v2) Antoun et al. (2020) model and for English we use bert-base-uncased model Devlin et al. (2019). Although metrics such as BLEU and ROUGE are commonly used, studies have reported their limitations Xu et al. (2023); Krishna et al. (2021). Therefore, we rely solely on BERTScore.
@@ 6 Experimental Results and Discussion
This section first presents competitive results among our proposed method and the state-of-the-art approaches. Next, it briefly analyzes and investigates the proposed method to validate and highlight the core contributions of this research.
Table 7 compares our proposed models with state-of-the-art approaches. On the ArMeme dataset, our method achieves the best accuracy at 72.1% and the best weighted F1 at 0.699, with Qarib and mBERT following behind. Although Qarib attains the highest macro F1 (0.551), our model remains competitive with a macro F1 of 0.536. Importantly, our method stands out because it provides explanations that add significant value. On the Hateful Meme dataset, our approach clearly outperforms the state-of-the-art by achieving the best performance with an accuracy of 79.9%, a weighted F1 of 0.802, and a macro F1 of 0.792. These results clearly highlight the advantages of our explainability-enhanced dataset and the proposed multi-stage optimization procedure for both classification and explanation-generation tasks.
@@
Table 7: Comparison with SOTA and our results. ResNet50 He et al. (2016) is an image only model. MS: Multi-stage.
Table 8 provides classification and explanation-generation results on the ArMeme and Hateful Meme datasets. It briefly presents these results from several perspectives: (a) Base vs. FT: demonstrates the performance difference between the same model with and without fine-tuning (FT); (b) Single-stage (SS) vs. Multi-stage (MS): highlights the necessity and benefits of the proposed optimization procedure and (c) Eng-Exp vs. Ar-Exp: showcases the multilingual capability of the selected VLM. Next, we provide a brief analysis of the results based on these perspectives.
First, we compare the Base vs. FT setup, from which it is evident that the FT model significantly outperforms the baseline. For example, on the ArMeme dataset, while the baseline achieves an accuracy of 12.7%, the proposed fine-tuning boosts it to 72.1%. Similarly, on the Hateful Meme dataset, fine-tuning improves the base accuracy from 65.2% to 79.9%. We observe similar improvements in the F1 metrics for classification and BERTScore for explanation quality. These significant performance gains validate our approach of fine-tuning the base models with the explainability enhanced dataset, demonstrating its efficacy for the meme classification and explanation generation tasks.
Next, we compare the SS vs. MS setup, which reveals that multi-stage (MS) fine-tuning further enhances performance over the single-stage (SS) approach. For example, on the ArMeme dataset, the accuracy increased from 68.2% to 72.1%, the weighted F1 increased from 0.584 to 0.699, the macro F1 increased significantly from 0.257 to 0.536, and the BERTScore for Arabic explanation increased significantly from 0.58 to 0.72. A similar trend is observed on the Hateful Meme dataset, where additional fine-tuning iterations yield more robust classification (approximately 4% improvement) and enhanced explanation quality. These performance gains validate our proposed multi-stage optimization procedure to further refine the VLMs.
Finally, we assess the model’s multilingual capability by comparing the performance of Llama MS - FT with Llama MS Ar-Exp. The results show that fine-tuning using explanations generated in both languages yields comparable outcomes. This validates the multilingual capability of our empirically chosen VLM for the target task and enables users to understand multilingual content even if they are not fluent in that language. For example, our model allows an English speaker to analyze Arabic memes and receive explanations in English.
Model Setup Acc (%) W-F1 M-F1 BS
@@
Llama Base 12.7 0.165 0.105 0.61
Llama SS FT 68.2 0.584 0.257 0.70
Llama MS FT 72.1 0.699 0.536 0.70
Llama Ar-Exp Base 19.0 0.246 0.125 0.58
Llama MS Ar-Exp FT 72.0 0.696 0.499 0.72
@@
Llama Base 65.2 0.615 0.567 0.661
Llama SS FT 75.9 0.760 0.745 0.767
Llama MS FT 79.9 0.802 0.792 0.777
Table 8: Results with ArMeme and Hateful meme classification and explanation generation. Llama: Llama-3.2 (11b), BS: BERTScore. SS: Single-stage, MS: Multi-stage. Ar-Exp: Model trained with Arabic explanation.
@@ 7 Conclusions and Future Work
In this study, we introduce a MemeXplain dataset for propagandistic and hateful meme detection and natural explanation generation, making it the first resource of its kind. To address both detection and explanation generation tasks and ensure efficient VLMs model training on this dataset, we also propose a multi-stage optimization procedure. To evaluate the multilingual capability of the model, we developed Arabic and English explanations for Arabic memes. The inclusion of English explanations benefits non-Arabic speakers, whereas providing explanations in the native language ensures that cultural nuances are accurately conveyed. With our multi-stage training procedure, we demonstrate improved detection performance for both ArMeme and hateful memes. The higher performance of explanation generation further demonstrates the efficacy of our multi-stage training approach. We foresee several future directions to extend this research and explore the following: (a) training the model with additional data through data augmentation, which could help it become an instruction-generalized model and potentially enhance its performance further; (b) incorporating pseudo and self-labeled data using an active learning procedure to incrementally improve the model’s capabilities; and (c) developing a task-generalized model that addresses multiple tasks.

### v2

@@ 4.3 Basic Statistics
Table 3 presents the basic statistics for both datasets. The average explanation length is 94 words for Arabic and 85 words for English. Notably, we instructed GPT-4o to generate explanations with fewer than 100 words.
@@ 5.1 Instructions Dataset
Our approach follows the standard pipeline for aligning LLMs with user intentions and specific tasks through fine-tuning on representative data Zhang et al. (2023); Kmainasi et al. (2024); Hasan et al. (2025). This process typically involves curating and constructing instruction datasets that guide the model’s behavior, ensuring it generates responses that align with the desired objectives. For our study, the responses include label and explanation. Hence, we created instruction format for both datasets. For the ArMeme dataset, we replicated the experiments for both Arabic and English explanations.
@@ 5.2 Model Selection
As shown in Figure 1, our first experimental phase involves model selection among several recent VLMs, including Llama-3.2 (11b) Dubey et al. (2024), Paligemma 2 (3b) Steiner et al. (2024), Qwen2-vl Wang et al. (2024), and Pixtral (12b) Agrawal et al. (2024). The goal here is to select the VLM that maximizes $p(l|i)$ for all image samples.
We evaluate the base models in a zero-shot setting and fine-tune them using an instruction-following paradigm. The instructions prompt the model to generate responses in the format ‘‘Label: (class_label)’’. We use a regex-based function to extract the predicted labels.
@@ 5.3 Multi-Stage (MS) Optimization Procedure
To emphasize our contribution, we introduce a novel optimization procedure to train VLM with MemeXplain, which decouples the classification and explanation generation tasks. While training both tasks in a single step may seem like an obvious choice, in practice, the two objectives produce conflicting gradient signals due to the fundamental differences between the learning objectives. Classification requires precise mapping of multimodal cues to discrete labels, whereas explanation generation demands fluency and coherence in free-form natural language. Merging them too early can compromise the model’s performance on both tasks. Conversely, training them completely separately risks overwriting knowledge of one task while learning the other. To address these challenges, we propose a two-stage optimization procedure that decouples the learning objectives by first optimizing in the classification domain in stage-1, followed by augmentation through explanation generation in stage-2. The overall goal of this optimization procedure is to obtain an optimal VLM that maximizes $p(l,e|i)$ for the training dataset (cf. Continual Training in Figure 1).
Problem Formulation.
We formulate the joint optimization problem as follows: $L_{\mathrm{total}}=L_{\mathrm{classif}}+W_{\mathrm{expl}}\;\cdot\;L_{\mathrm{expl}},$ where $L_{\mathrm{classif}}$ and $L_{\mathrm{expl}}$ are the classification and explanation losses, respectively. $W_{\mathrm{exp}}$ is a step‑function weight that switches from 0 to 1 between the stages. This formulation draws on the principles from Domain Adaptation and Task‑Incremental Learning (Van de Ven et al., 2022) with the aim to: (a) Isolate classification learning and avoid conflicting updates and (b) Prevent catastrophic forgetting by gradually integrating the explanation objective.
Stage 1 - Classification Fine-Tuning.
We set $W_{\mathrm{exp}}=0$ and optimize exclusively on $L_{\mathrm{classif}}$. It adapts the pretrained VLM to the domain of hateful/propagandistic content, establishing a strong feature backbone for accurate label prediction.
Stage 2 - Joint Classification & Explanation Fine‑Tuning.
We set $W_{\mathrm{exp}}=1$ to optimize $L_{\mathrm{classif}}+L_{\mathrm{explanation}}$. After obtaining the domain-adapted backbone from Stage 1, the model proceeds to learn how to generate coherent, contextually grounded explanations alongside accurate classifications. This stepwise integration ensures that the model preserves its classification capabilities and avoids catastrophic forgetting, while developing proficiency in natural language reasoning.
By decoupling and then recombining these objectives in a straightforward two-phase procedure, our multi-stage (MS) optimization presents an easily implementable extension to standard VLM training pipelines. To validate its effectiveness, we compare it against a single-stage (SS) fine-tuning baseline, where the model is directly trained on the label-with-explanation dataset. Our ablation studies (detailed in Section 6) demonstrate that the proposed multi-stage approach significantly outperforms the single-stage strategy.
@@
Qwen2V (7b) Base 63.1 0.550 0.242
Qwen2V (7b) FT 72.2 0.686 0.440
@@
Qwen2V (7b) Base 66.4 0.669 0.442
Qwen2V (7b) FT 77.9 0.773 0.753
@@ 5.4 Training Setup
Our fine-tuning experiments utilize QLoRA Dettmers et al. (2023), which combines INT4 quantization with parameter-efficient fine-tuning through Low-Rank Adaptation (LoRA) Hu et al. (2022). In our setup, the base model is quantized to 4-bit precision, with LoRA updates applied to a subset of the model parameters. This approach was selected to address computational and memory resource constraints. We adapted all relevant submodules (vision, language, attention, and MLP layers) with a LoRA rank of 16, an alpha of 16, and no dropout. We used a per-device batch size of 2 with gradient accumulation over 4 steps, optimizing with AdamW, a weight decay of 0.01, and a linear scheduler with 5 warmup steps.
Single-Stage (SS) Baseline.
We implement a baseline where models are fine-tuned end-to-end using a unified classification and classification-with-explanation objectives. Model selection is performed on the development set, and final results are reported on the test set using the best-performing checkpoint.
Multi-Stage (MS) Curriculum.
We design a staged curriculum to progressively instill multimodal reasoning capabilities into the model. In the first stage, the model is trained exclusively on the classification dataset, and intermediate checkpoints are evaluated on the development set. The best checkpoint from this stage is then used to initialize the second stage, where the model is further optimized with the classification-with-explanation objective. Hyperparameter exploration revealed that lowering the learning rate while increasing the number of training epochs in the second stage yields performance gains in our setup.
@@ 5.5 Evaluation Setup and Metrics
The results for pretrained models, fine-tuned models, and baselines are summarized in Tables 5 and 6. For performance measurement across different experimental settings, we compute accuracy, weighted F1, and macro-F1. We evaluate the model’s explanation performance on the test set using semantic similarity-based metric, measured by the F1 score within BERTScore (Zhang et al., 2020). This score is computed using contextual embeddings extracted from pre-trained BERT models. To enhance accuracy, we utilize language-specific transformer models for embedding extraction. For Arabic we use AraBERT (v2) Antoun et al. (2020) model and for English we use bert-base-uncased model Devlin et al. (2019). In addition, we also compute BLEU and METEOR scores.
@@ 6 Experimental Results and Discussion
This section first presents competitive results among our proposed method and the state-of-the-art approaches. Next, it briefly analyzes and investigates the proposed method to validate and highlight the core contributions of this research. To increase clarity, we emphasize that all the results discussed in this section are based on an experimental setup that evaluates models using the MemeXplain dataset, which includes both classification and explanation for training. This differs from the results presented in the previous section, in Tables 5 and 6, where we evaluated and compared models trained on the original datasets, which include only classification labels. Therefore, none of the results discussed in this section should be compared to those presented in Tables 5 and 6.
Comparison with State‑of‑the‑Art.
Table 7 presents the comparison. On the ArMeme dataset, our proposed approach (Llama MS) achieves the highest accuracy at 72.1% and the best weighted F1 score at 0.699, with Qarib and mBERT trailing behind. Although Qarib achieves the highest macro F1 score (0.551), Llama MS remains competitive with a macro F1 of 0.536. Importantly, our approach stands out by providing explanations that add significant value. On the Hateful Meme dataset, Llama MS surpasses the state-of-the-art by achieving the best performance: an accuracy of 79.9%, a weighted F1 score of 0.802, and a macro F1 score of 0.792. These results clearly highlight the advantages of our explainability-enhanced dataset and the effectiveness of our proposed optimization procedure for both classification and explanation-generation tasks.
@@
Table 7: Comparison with SOTA and our results (Llama MS and Llama (Ar-Exp) MS). ResNet50 He et al. (2016) is an image only model. MS: Multi-stage (Our approach).
Model Setup Acc (%) W-F1 M-F1 BS BL M
@@
Llama Base 12.7 0.165 0.105 0.61 0.24 0.17
Llama SS FT 68.2 0.584 0.257 0.70 0.56 0.36
Llama MS FT 72.1 0.699 0.536 0.70 0.57 0.35
Llama Ar-Exp Base 19.0 0.246 0.125 0.58 0.12 0.09
Llama MS Ar-Exp FT 72.0 0.696 0.499 0.72 0.55 0.29
@@
Llama Base 65.2 0.615 0.567 0.661 0.35 0.23
Llama SS FT 75.9 0.760 0.745 0.767 0.65 0.47
Llama MS FT 79.9 0.802 0.792 0.777 0.67 0.49
Table 8: Results with ArMeme and Hateful meme classification and explanation generation. Llama: Llama-3.2 (11b), BS: BERTScore, BL: BLEU, M: METEOR. $S$: Single-stage, MS: Multi-stage. Ar-Exp: Model trained with Arabic explanation. W-F1: Weighted F1, M-F1: macro-F1
Ablation Study on Different Model Training Settings.
Table 8 presents classification and explanation-generation results on the ArMeme and Hateful Meme datasets across different settings: (a) Llama represents the performance of the base model without fine-tuning; (b) Llama SS represents the performance after single-stage fine-tuning; and (c) Llama MS represents the performance after completing the second stage of our proposed multi-stage optimization. The extension Ar-Exp after a setting name indicates that the model was trained using Arabic explanations. Based on these settings, we analyze the results from various perspectives: (a) Base vs. FT: demonstrates the performance difference between the same model with and without fine-tuning (FT);
(b) Single-stage (SS) vs. Multi-stage (MS): highlights the necessity and benefits of the proposed optimization procedure; and
(c) Eng-Exp vs. Ar-Exp: showcases the multilingual capability of the selected VLM.
First, we compare the Base vs. FT setup in Table 8, from which it is evident that the FT model significantly outperforms the baseline. For example, on the ArMeme dataset, while the baseline (Llama) achieves an accuracy of 12.7%, the proposed fine-tuning (Llama-MS) boosts it to 72.1%. Similarly, on the Hateful Meme dataset, fine-tuning improves the base accuracy from 65.2% (Llama) to 79.9% (Llama-MS). We observe similar improvements in the F1 metrics for classification and BERTScore for explanation quality. These significant performance gains validate our approach of fine-tuning the base models with the explainability enhanced dataset, demonstrating its efficacy for the meme classification and explanation generation tasks.
Next, we compare the SS vs. MS setup in Table 8, which reveals that multi-stage (MS) fine-tuning further enhances performance over the classification single-stage approach. For example, on the ArMeme dataset, the accuracy increased from 68.2% (Llama-SS) to 72.1% (Llama-MS), the weighted F1 increased from 0.584 (Llama-SS) to 0.699 (Llama-MS), the macro F1 increased significantly from 0.257 (Llama-SS) to 0.536 (Llama-MS), and the BERTScore for Arabic explanation increased significantly from 0.58 (Llama-SS) to 0.72 (Llama-MS). A similar trend is observed on the Hateful Meme dataset, where additional fine-tuning iterations yield more robust classification (approximately 4% improvement) and enhanced explanation quality. These performance gains validate our proposed multi-stage optimization procedure to further refine the VLMs.
Assessment of Multilingual Capability.
We compare Llama MS - FT with Llama MS Ar-Exp - FT in Table 8, which shows that fine-tuning using explanations generated in both languages yields comparable outcomes. This validates the multilingual capability of our empirically chosen VLM for the target task and helps users understand multilingual content without fluency in that language. For example, our model allows an English speaker to analyze Arabic memes and receive explanations in English.
Annotation Agreement: Additional Analysis
We report human evaluations of the generated explanations from two models: (i) Llama MS (ArMeme) and (ii) Llama MS (Hateful Memes), as detailed in Table 9. Due to resource constraints, the manual evaluation was conducted on a representative sample of 100 instances. For the ArMeme dataset, the scores across all evaluation metrics range from 4.15 to 4.74 out of 5, while for the Hateful Memes dataset, the scores range from 4.41 to 4.54, indicating consistently strong performance across both datasets.
Dataset Faithfulness Clarity Plausiblity Informative
ArMeme 4.63 4.74 4.56 4.15
Hateful meme 4.41 4.44 4.43 4.54
Table 9: Average Likert scale value for each human evaluation (annotation) metric for the explanations of different datasets.
Error Analysis.
To better understand the model’s reasoning capabilities and failure modes, we conduct an error analysis contrasting correct and incorrect label predictions, as well as comparing single-stage and multistage training paradigms. Our analysis reveals that while the model often aligns well with human explanations in correctly classified instances, it struggles with nuanced cases involving humor or implicit context. Furthermore, multistage training consistently outperforms single-stage training in label accuracy by better integrating textual and visual cues while providing explanation. See Section D for detailed examples and discussion.
@@ 7 Conclusions and Future Work
In this study, we introduce a MemeXplain dataset for propagandistic and hateful meme detection and natural explanation generation, making it the first resource of its kind. To address both detection and explanation generation tasks and ensure efficient VLMs model training on this dataset, we also propose a multi-stage optimization procedure. To evaluate the multilingual capability of the model, we developed Arabic and English explanations for Arabic memes. The inclusion of English explanations benefits non-Arabic speakers, whereas providing explanations in the native language ensures that cultural nuances are accurately conveyed. With our training procedure, we demonstrate improved detection performance for both ArMeme and hateful memes. The higher performance of explanation generation further demonstrates the efficacy of our multi-stage training approach.
Moreover, the proposed multi-stage scheme is agnostic to specific VLM architectures and applicable in contexts where tasks impose divergent gradient demands, thereby offering a versatile framework for multi-task vision–language learning.
We foresee several future directions to extend this research and explore the following: (a) training the model with additional data through data augmentation, which could help it become an instruction-generalized model and potentially enhance its performance further; (b) incorporating pseudo and self-labeled data using an active learning procedure to incrementally improve the model’s capabilities; and (c) developing a task-generalized model that addresses multiple tasks.
@@ Appendix D Error Analysis
D.2 Effect of Single vs. Multistage Training
In Figure 4 we notice that multi-stage model correctly predicts the label and generates a faithful explanation, identifying how the combination of the image and text reinforces gender stereotypes. In contrast, the single-stage model label the meme with incorrect label – not-propaganda. Figure 5 shows another example from the hateful memes dataset. The multi-stage model correctly labels the image as hateful and grounds its explanation in the offensive combination of the swastika and the mocking caption. Whereas the single-stage model misclassifies the image as not-hateful. This highlights the effectiveness of multistage optimization in aligning predictions with grounded reasoning.
@@
Figure 4: Single- vs. Multi-Stage prediction for ArMeme.
Figure 5: Single- vs. Multi-Stage prediction for hateful memes.